PLEASE please be very very clear and please show me step by step please2. Consider the dynamical system x = I – ry y + 2xy y Find the equilibrium points, and draw the isoclines of horizontal and
Answer #1

ANSWER:

Solution : consider the dynamic System x = x-2y y = y +224 Approximate linear form is : f(x,y) ~ F(0.0) +2 90+(3:1). 0+(l) +2 Matrix we be z -2 2 values are x²- 22+G4)=0 eigen x²-22-4=0 pe At x=0 > y() 200 at fixed constant at. X=0 equilibrium state